OpenStudy (anonymous):

Let F(x) = 4/x. Using the definition of a derivative, write f'(2) as a limit in two different ways. f'(2)=lim h->0 numerator1/h where numerator1 = ? and f'(2)=lim x->2 numerator2/x-2 where numerator2 = ? We will find shortcuts for computing derivatives in Chapter 3. For now, it is good practice to compute the limit above using methods of section 2.3. Using your choice of either limit above, compute . f'(2) Answer : ? Working on this for a while. Can someone help me?

OpenStudy (tkhunny):

\(\dfrac{f(x+h) - f(x)}{h}\) \(\dfrac{\dfrac{1}{x+h} - \dfrac{1}{x}}{h}\)

OpenStudy (tkhunny):

* 4, not 1

OpenStudy (anonymous):

it says the answer is wrong

OpenStudy (tkhunny):

Sadly, you provided nothing of your work. This makes it VERY difficult to know what you did and how you presented what answer that was wrong. Please do better at communication.